Identify the solution(s) of the system of equations, if any.
6x + 4y = 8
y = 5 - 3x

Answer by jim_thompson5910(35256)

Start with the first equation and plug in y = 5 - 3x. Then solve for x.


6x + 4y = 8


6x + 4(5 - 3x) = 8


6x + 20 - 12x = 8


-6x + 20 = 8


-6x = 8 - 20


-6x = -12


x = -12/(-6)


x = 2


-------------------


Now that you know that x = 2, plug it into y = 5 - 3x to find y.


y = 5 - 3x


y = 5 - 3(2)


y = 5 - 6


y = -1


-------------------


Answer: x = 2, y = -1 giving the ordered pair (2, -1)